log9 (x-3) + log9 (x-3)= 1
Answer by fractalier(2101)

Okay from
log9 (x-3) + log9 (x-3)= 1
we combine to get
2*log9 (x-3) = 1
log9 [(x-3)^2] = 1
now exponentiate "nine to the..." and get
[(x-3)^2] = 9
x - 3 = ±3
and x = 6 or x = 0
but x cannot be zero, so
x = 6